ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 206, 2015

Adrian, M. and J. Lansky An interpretation of the tame local Langlands correspondence for p-adic P GSp(4) from the perspective of real groups

353

Aizenbud, A., D. Gourevitch and S. Sahi Derivatives for smooth representations of GL(n, ℝ) and GL(n, ℂ)

1

Aizenbud, A., D. Gourevitch and S. Sahi Twisted homology for the mirabolic nilradical

39

Asensio, P. A. G., D. K. Tütüncü and A. K. Srivastava Modules invariant under automorphisms of their covers and envelopes

457

Bárány, I. and V. S. Grinberg Block partitions of sequences  169

Bergen, J. and P. Grzeszczuk Gelfand–Kirillov dimension of algebras with locally nilpotent derivations

313

Bertapelle, A. and C. D. González-Avilés On the cohomology of tori over local fields with perfect residue field

431

Bierstone, E. and F. V. Pacheco Desingularization preserving stable simple normal crossings

247

Bourgain, J. Prescribing the binary digits of primes, II  179

Brudnyi, A. A Rolle Type Theorem for Cyclicity of Zeros of Families of Analytic Functions

95

Chorny, B. Homotopy theory of relative simplicial presheaves  109

Damanik, D. and A. Gorodetski Almost ballistic transport for the weakly coupled Fibonacci Hamiltonian

123

Eick, B. and D. J. Green The Quillen categories of p-groups and coclass theory

197

Emamizadeh, B. and Y. Liu Constrained and unconstrained rearrangement minimization problems related to the p-Laplace operator

295

García-García, J. I., M. A. Moreno-Frías and A. Vigneron-Tenorio Computation of the ω-primality and asymptotic ωprimality with applications to numerical semigroups

395

González-Avilés, C. D. See A. Bertapelle Gorodetski, A. See D. Damanik Gorodski, C. and F. Podestà Tight Lagrangian homology spheres

in compact homogeneous Kähler manifolds

413

Gourevitch, D. See A. Aizenbud Green, D. J. See B. Eick Grinberg, V. S. See I. Bárány Grzeszczuk, P. See J. Bergen Guo, Y. and G. Wei Determination of differential pencils from dense

nodal subset in an interior subinterval

227

Lansky, J. See M. Adrian

Lewis, M. L. Camina pairs that are not p-closed  89

Liu, Y. See B. Emamizadeh Malicki, M. Generic elements in isometry groups of Polish ultramet-

ric spaces

141

Mattarei, S. Inversion and subspaces of a finite field  327

Moreno-Frías, M. A. See J. I. García-García Pacheco, F. V. See E. Bierstone Podestà, F. See C. Gorodski Sahi, S. See A. Aizenbud Srivastava, A. K. See P. A. G. Asensio Tütüncü, D. K. See P. A. G. Asensio Tamuz, O. and R. J. Tessler Majority Dynamics and the Retention

of Information

483

Tessler, R. J. See O. Tamuz Vigneron-Tenorio, A. See J. I. García-García Wei, G. See Y. Guo